OpenStudy (pinkcandyrosez882):

For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction If you can, please explain your answer and why! Thanks! (Will give medal!)

OpenStudy (pinkcandyrosez882):

@Thesmarterone @texaschic101 @nevermind_justschool @Angelina14

OpenStudy (jango_in_dtown):

A. multiplication. see 1*0=0 0*1=0 1*1=1 0*0=0 so when we multiply any two members of the set, it gives the member of the same set

OpenStudy (jango_in_dtown):

division by 0 is not possible. so option B is not true

OpenStudy (pinkcandyrosez882):

Thanks so much

OpenStudy (jango_in_dtown):

also see that 1+1=2 which is not a member of the set. hence option c is also not true

OpenStudy (jango_in_dtown):

finally 0-1=-1 which is not a member of the set. So option d is also not true. and hence there is only one correct option which is option A

OpenStudy (pinkcandyrosez882):

Thanks! This helps! :)